OpenStudy (anonymous):

At 1 pm ship A is 5 km due south of ship B. Ship A sails north at 3 km/h and ship B sails west at 4 km/h. When is the distance between the ships a minimum? how far apart are they at this time.

OpenStudy (anonymous):

d² = ( 5 - 3t)² + (4t )² = 25t² - 30t + 25 ->2dd' = 50t - 30 => d' = 0 -> t = 3/5 hr Thus Min d = √ ( 5 - 3* 3/5 )² + (4 * 3/5 )² = ...